Grade 11Algebra

if the odds in favour of an event be 3:5, then find the probability of non-occurrence of the event is

Sakkshi
8 Years agoGrade 11

3 Answers

nikhil kr
8 Years ago
Total Probability=1 So Probability of Occurance+Probability of non occurance=1 Given that Probability of occurance of the event is (3/5) So 3/5+Probablity of non occurance=1 Probability of non occurance=1-(3/5) = (5-3)/5= (2/5)So the Probability of non occurance of the given event is 2/5
Arun
8 Years ago
Dear student
 
Odds in favour = 3:5
Hence
Probability of occurance = 3/8
Hencw probability vof non occurance = 1 - 3/8 = 5/8
Rishi Sharma
5 Years ago
Dear Student,
Please find below the solution to your problem.

Let E be an event
Number of occurrence events =3
Number of non-occurrence events =5
Total number of event =3+5=8
Probability of occurrence = 3/8
Hencw probability of non occurance = 1 - 3/8 = 5/8


Thanks and Regards